Grade 11Mechanics

A car is to be hoisted by elevator to the fourth floor of a parking garage, which is 14 m above the ground. If the elevator can have a maximum acceleration of 0.2 m/s^(2) and maximum decelaration of 0.1m/s^(2) and can rach a maximum speed of 2.5 m/s, determine the shortest time to make the lift, starting from rest and ending at rest.

Pritesh Ramya
10 Years agoGrade 11

3 Answers

Vikas TU
9 Years ago
USE third law of eqn.
v^2 = u^2 + 2as
for 
2.5^2 = 0^2 + 2*0.2*s
find s from here then compare with 14 meter.
dgsgadsgfdahfda
8 Years ago
hehehehehemhgtkgkhlkjh;ljhuojholjholhlhiklhlikhlihlihlihlik;hkip'gkp'gfljf
[pl6j’;lk;jhjlk;'gdkhpogfdkhtr
 gergtfskh;'trs khtrku[s’6ypea’rkhp;'atrk
uhyean,fds/mkmrm;a/.jbv
dgsgadsgfdahfda
8 Years ago
fuck you nobe!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!why do you go to school??????????????????????????????????????????????????????